OpenStudy (anonymous):

A cone has a base that is 12 1/2 cm2 and a height of 10 cm. what is the volume

Mehek (mehek14):

do you know the formula?

Mehek (mehek14):

\(\large\dfrac{πr^2h}{3}\)

OpenStudy (anonymous):

but i don't know the radius

Mehek (mehek14):

your welcome

Mehek (mehek14):

do you by chance have any answer choices?

OpenStudy (anonymous):

no i do not

Mehek (mehek14):

ok now the base is a circle the formula for that is \(\large{πr^2}\)

Mehek (mehek14):

the total base is 12.5 so divide 12.5/3.14

OpenStudy (anonymous):

39.25

OpenStudy (anonymous):

thanks

Mehek (mehek14):

12.5/3.14 = 3.98 now find the \(\large\sqrt{3.98} = 1.99~which~can~be~rounded~to~2 \)

Mehek (mehek14):

now use the formula for the volume of the cone